Find the characteristic values and the characteristic functions of the sturn_ liouville problem.

d/dx[x(dy/dx)]+(λ/x)y=0

y1(1)=0,y1(e2x)=0 where λ is nonnegative


1
Expert's answer
2021-10-26T09:38:08-0400

if "\lambda=0" :

"xy'=c_1"

"y=c_1lnx+c_2"

"y'(1)=c_1=0"

"y=c_1"


if "\lambda>0" :

"\lambda=k^2" with "k>0"

"x^2y''+xy'+k^2y=0"

an Euler equation with indicial equation:

"r^2+k^2=(r-ik)(r+ik)=0"

"y=c_1cos(klnx)+c_2sin(klnx)"


"y'=-c_1sin(klnx)\/x+c_2cos(klnx)\/x"

"y'(1)=c_2=0"

"y'(e^{2x})=-c_1sin(2kx)\/e^{2x}=0"

"2kx=\pi n"

"k=\pi n/(2x)"


"\lambda_n=(\pi n/(2x))^2"

"y_n=cos(\pi nlnx/2x)"